OpenStudy (anonymous):

how do i solve g(x)=-3x2) (the last bit is 3x squared

OpenStudy (unklerhaukus):

but it is solved

OpenStudy (anonymous):

no i have to solve it with the quadratic function y=a(x-h)2+k (once again the 2 means squared

OpenStudy (unklerhaukus):

so g(x)=y? \[-3x^2 = a(x-h)^2+k\]

OpenStudy (unklerhaukus):

a=-3, h=0, k=0